# Observation of the quantum equivalence principle for matter-waves

### v3

@@ VII.7.2 Derivation of the phase difference using the action
which, in this case, gives
$a_{kick}\cdot 2T_{kick}=a_{hold}\cdot T_{h}\,.$ (16)
Setting $a_{hold}=g$ confirms that the reference path is stationary in the lab frame. Performing the integrals, we get
$\phi_{1}=\frac{mg^{2}}{3\hbar}\left(-T^{3}-T^{2}T_{kick}-TT_{kick}^{2}-TT_{kick}T_{d}+2T_{kick}^{3}+7T_{kick}^{2}T_{d}+8T_{kick}T_{d}^{2}+3T_{d}^{3}\right)$ (17)
$\phi_{2}=\frac{2mg^{2}}{3\hbar}\left(T_{kick}+T_{d}\right)^{3}$ (18)
$\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}+T^{2}T_{kick}+T\left(T_{kick}^{2}+T_{kick}T_{d}\right)-T_{d}\left(T_{kick}+T_{d}\right)^{2}\right).$ (19)
Taking the limit of a short kick we get
$\lim_{T_{kick}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}-T_{d}^{3}\right),$ (20)
and the limit of short delay we get
$\lim_{T_{kick},T_{d}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}T^{3}.$ (21)
Eq.21 represents the Gedanken experiment in which the beam splitter is a delta pulse such that $T_{kick},T_{d}\to 0$.
The result of the analytical model given in Eq. (19) (and in Eq. (4) of the main text) is plotted in Fig. 3 of the main text using the experimental values of $T_{kick}=80\rm\,\mu s$ and $T_{d}=71+6\rm\,\mu s$ where we include half the rise time of the holding pulse $\tau_{\rm hold}/2=6\rm\,\mu s$ in the delay. We also include the contribution to the phase from the force on the state $|0\rangle$ due to the second-order Zeeman (SOZ) effect. This force is opposite to the force on the $|1\rangle$ state during the gradient pulses, and its value is about 1% of the latter. During the holding pulse, the acceleration of the state $|0\rangle$ due to SOZ is calculated to be ${a}_{SOZ}=0.10\,\rm m/s^{2}$ (see Sec. K) for details). We include this effect in the analytical prediction for the phase (Eq. 17) by replacing the gravitational acceleration $g$ by $g_{\rm eff}=g+a_{\rm SOZ}=9.91\,\rm m/s^{2}$. This estimation agrees with a model in which $a_{SOZ}$ is explicitly included in the calculation to within $0.03\,\rm rad$.
VII.7.3 Calculating the actions in both frames
Next, to understand the symmetry of the QGI, as emphasized in the introduction, we calculate for the Gedanken experiment the actions associated with the classical trajectories of the reference and the ballistic wave packets as seen from both frames of reference, keeping track of the contributions from the kinetic and potential energies as well as the pulses, which we model for the sake of simplicity by delta functions.
We start by discussing the situation in the Newtonian frame. Here the reference wave packet remains at rest for all times, since the gravitational acceleration $g$ is compensated by the acceleration $a$ provided by the magnetic field gradient. Hence, the contributions due to the kinetic and potential energies vanish. Moreover, there are no pulses. Thus, the total action of the reference wave packet vanishes as indicated in Table 1.
In contrast, two parts contribute to the total action of the ballistic wave packet. Due to the non-vanishing momentum $mv_{0}$ of the atom originating from the first pulse, the accelerated motion in the linear gravitational potential provides us with non-vanishing actions due to the kinetic as well as the potential energies. Their structure shown in the second row of Table 1 becomes apparent when we recall that the velocity of this motion consists of the constant velocity $v_{0}$ and the one caused by the gravitational acceleration $g$. The action due to this kinetic energy involves the integral of the square of this time-dependent velocity creating the sum of three terms: ($i$) The kinetic energy $mv_{0}^{2}/2$ multiplied by the interferometer time $2T$. ($ii$) A cross term between the velocity $v_{0}$ and the acceleration $g$ and ($iii$) the square of the acceleration times $(2T)^{3}$.
This scaling in $v_{0}g$ and $g^{2}$ is crucial for the equivalence principle. Indeed, when we replace $v_{0}$ by $-v_{0}$ and $g$ by $-a$ as defined by the Einsteinian frame discussed in the second column of Table 1, the expressions, now occurring for the reference rather than the ballistic wave packet, are identical.
Next we turn to the contribution from the potential energy of the ballistic wave packet in the Newtonian frame. Since the ballistic wave packet starts from the origin of the coordinate axis, the motion is a combination of the linear motion due to $v_{0}$, and the gravitational acceleration $g$ giving rise to two terms in the action which are identical in size and sign to the last two terms of the action due to the kinetic energy. Hence, these terms also satisfy the symmetry in $v_{0}g$ and $g^{2}$.
It is interesting that in the Newtonian frame the delta function pulses do not contribute to the actions of the ballistic wave packet in the Newtonian frame, and this for two reasons: First, the corresponding potential energy is proportional to the coordinate at the time of the pulse and second, the ballistic wave packet starts at $z=0$ and concludes its journey at $z=0$.
As a result, the total action of the ballistic wave packet is given by the ones of the kinetic and potential energies which due to the closing condition connecting $v_{0}$ and $T$ allows us to combine the three terms in a single one shown in the Table 1, determining the difference of the actions corresponding to the two wave packets.
Finally, we address the Einsteinian frame which is attached to the motion of the ballistic wave packet. As mentioned before, the two wave packets exchange their roles compared to the Newtonian frame. Hence, now the action of the ballistic wave packet vanishes but the one of the reference wave packet is non-vanishing arising from the magnetic field gradient in the same way as for the ballistic wave packet in the Newtonian frame.
Moreover, the symmetry of the terms discussed before, ensures that the origin of this contribution is identical to one of the terms of the ballistic wave packet in the Newtonian frame.
Since we consider the difference of the actions of the two wave packets there must be a sign difference between the two frames. Indeed, in the Newtonian frame the action of the reference wave packet vanishes, whereas for the Einsteinian frame it is that of the ballistic wave packet.
Table 1: Contributions to the actions of the center-of-mass motion of the reference and the ballistic wave packets in the QGI viewed from the Newtonian and the Einsteinian frames. Note that the closing conditions to obtain the respective total actions are written in round brackets whereas the levitation condition for the respective difference in actions are given by the expression in square brackets. In the Newtonian frame the reference wave packet is levitated and remains at rest at the origin. The Einsteinian frame is attached to the ballistic wave packet. Thus, in this frame the wave packet remains at rest at the origin yielding only vanishing actions. This is also the case for the reference wave packet but in the Newtonian frame.
actions frame Newtonian Einsteinian
reference wave packet kinetic 0 $mv_{0}^{2}T-2mv_{0}aT^{2}+\frac{4}{3}ma^{2}T^{3}$
potential 0 $-2mv_{0}aT^{2}+\frac{4}{3}ma^{2}T^{3}$
pulses no pulses 0
total 0 $-\frac{1}{3}ma^{2}T^{3}$ $(v_{0}=aT)$
ballistic wave packet kinetic $mv_{0}^{2}T-2mv_{0}gT^{2}+\frac{4}{3}mg^{2}T^{3}$ 0
potential $-2mv_{0}gT^{2}+\frac{4}{3}mg^{2}T^{3}$ 0
pulses 0 no pulses
total $-\frac{1}{3}mg^{2}T^{3}$ ($v_{0}=gT$) 0
difference in total actions $-\frac{1}{3}mg^{2}T^{3}$ $\frac{1}{3}ma^{2}T^{3}$ $[a=-g]$
VII.7.4 Derivation of the phase difference using Galilean transformation
Finally, we rederive the phase difference measured by the QGI using the combination of the Galilean transformation and the one into freely-falling frame, given by Eq. (1) of the main text. For this purpose, we first recall the Galilei transformation [63]
$\psi_{N}(z,t)=e^{i\varphi(z,t)}\psi_{0}\left(z-v_{0}t,t\right),\;\text{where}\;\;\;\varphi(z,t)\equiv\frac{m}{\hbar}\left(-\frac{1}{2}v_{0}^{2}t+v_{0}z\right)$ (22)
between the time-dependent wave functions in the Newtonian frame and in the frame moving with the constant velocity $v_{0}$. Due to the momentum transfer $p_{0}=mv_{0}$ onto the initial wave packet $\psi_{0}(z,0)$, the particle is at rest in this frame but undergoes free quantum dynamics.
To include the gravitational acceleration, we combine Eq. (1) and Eq. (20) which yields
$\psi_{N}(z,t)=e^{i\phi_{gauge}(z,t)}e^{i\varphi\left(z+\frac{1}{2}gt^{2},t\right)}\psi_{0}\left(z-v_{0}t+\frac{1}{2}gt^{2},t\right).$ (23)
We now consider the wave function $\psi_{N}$ at $t=2T$ and use the closing condition $T=v_{0}/g$. As a result, the spatial argument in $\psi_{0}$ reduces to $z$. Moreover, the second pulse employed at $t=2T$ imprints again the momentum $p_{0}$ onto the wave function, resulting in
$\psi_{N}(z,2T)=e^{i\phi_{ballistic}}\psi_{0}\left(z,2T\right),$ (24)
where we have introduced the ballistic phase
$\phi_{ballistic}=-\frac{1}{3\hbar}mg^{2}T^{3}.$ (25)
VII.8 Numerical simulation
VII.8.1 Details of the numerical simulation
The numerical simulation calculates the dynamics of the atomic BEC during the trapping, release, DKC, and interferometer sequence. The simulation is based on the model of wave packet evolution given in [54], which is extended to include the effects of rotations. To use the model with curved magnetic potentials, we locally approximate the magnetic potential by a quadratic potential and calculate the evolution of each wave packet in the potential. The simulation predicts the output of the SGI by calculating the overlap integral of the two final wave packets, resulting in a complex value, where the visibility is the magnitude of this integral and the phase is the argument.
The simulation is constructed to mimic the exact experimental sequence by using the same timings as in the experiment. During the preparation stage, the currents and magnetic fields in the simulation are kept close to the values reported in the experiment but fine-tuned to match the wave packet’s initial conditions, such as initial position and velocity, and initial expansion rate. As the measurements of the absolute positions of the atoms are accurate up to a few $\rm\mu m$, other indications of the position of the atoms in the lower interferometer arm are also taken from the holding current required to hold the atoms against gravity, and the trajectory after the DKC. A holding current of 23-24 mA corresponds, according to the calculation, to a distance of 112-113 $\mu$m from the chip surface. Assuming that the velocity of the atoms at this position before turning on the holding pulse is almost zero, we can deduce the trajectory after the DKC as the atoms are in free fall and hence travel a distance $\Delta z=\frac{1}{2}gt^{2}$. This calculation indicates a distance of about 132 $\mu$m from the chip surface for the initial trapping position. These values of the initial trap and turning point agree with direct measurements from the experiment.
The magnetic bias field is taken to be the field generated by the bias coils and Earth’s magnetic field ${\bf B}_{E}=(-0.31,0,0.33)$ G. The bias generated by the coils consists of a component in the $y$ direction and in the $x$ direction, which are determined to be $B_{0y}=12.9928$ G and $B_{0x}=-1.175$ G. The $y$-component of the field during the trapping determines the trap position with respect to the chip surface and the $x$-component determines the trap frequency and field at the bottom of the trap. The total magnetic field magnitude is indeed measured some time after trap release by the RF transition frequency between the Zeeman levels $|2,2\rangle$. and $|2,1\rangle$. This measurement indicates that the total field at this time is $B_{0}=12.625$ G, but we take the field during the trapping to be $B_{0}=13.05$ G to ensure that the trapping position is about 132 $\mu$m from the chip surface, which is consistent with the position of the atoms in the lower arm of the interferometer, as described above. The discrepancy between the measured bias field after the trap release and the predicted bias field required to explain the trapping stage may be attributed to the magnetic response of the atom chip mount and the chamber itself, resulting in transient fields in the science chamber during the trapping. Time dependence of the bias field after the trap release and DKC is observed in the RF resonance measurements. The $z$-component of the magnetic field generated by Earth is responsible for a shift of the trapping position relative to the center of the system and we find numerically that the trapping position is at ${\bf r}_{\rm trap}=(0.79,4.67,132.05)\,\mu$m relative to the center of the surface of the central wire. The trapping frequencies are $\omega_{y}\approx\omega_{z}\approx 2\pi\times 1.04$ kHz and $\omega_{x}\approx 2\pi\times 57$ Hz, which is in good agreement with the measured trap frequencies. The calculated field at the bottom of the trap is $B_{0}\approx 0.825$ G, which is slightly larger than the measured value. For a total number of $2\cdot 10^{4}$ atoms, we obtain a BEC of an average width $\sigma_{x}=3.13\,\mu$m in the longitudinal direction and $\sigma_{y}=\sigma_{z}=1.31\,\mu$m in the transverse direction. This serves as the starting point for calculating the evolution of the wave packet widths during the sequence. The constant bias field also includes a measured gradient $B^{\prime}=68.32$ G/m, giving rise to an acceleration $\mu_{B}B^{\prime}/2m=0.22$ m/s2 in the direction of gravity. The DKC pulse kicks the atoms upwards, and after it, the atoms are in free fall until the beginning of the interferometer, except that they are affected by small magnetic gradients from the bias field and an idle current from the kick wires on the chip. During this time, they are transferred from the state $|2,2\rangle$ to the state $|2,1\rangle$ by a RF pulse that ends about 400 $\mu$s after the collimation pulse and then they are split into a superposition of $|2,1\rangle$ and $|1,0\rangle$. As the idle current is designed to counteract the magnetic gradient of the bias during the time between the splitting $\pi/2$ pulse and the recombining $\pi/2$ pulse we may consider the trajectory of the atoms before the kick pulse as being ballistic and calculate the timing and position of the turning point where the atoms are expected to have a zero velocity. This point $z_{h}$ is located at a distance $z_{h}-z_{\rm DKC}=\frac{1}{2}gt_{\rm free}^{2}$ from the position $z_{\rm DKC}$ at the end of the collimation pulse. At this point, the holding current $I_{h}$ is calculated and the kick current $I_{k}$ is also calculated such that the holding current and kick currents satisfy $2I_{k}T_{k}=I_{h}T_{h}$. The DKC is optimized to yield a zero velocity of the atoms at the beginning of the holding pulse such that the lower trajectory is constant during the holding time. The ballistic trajectory is calculated based on the calculated kick current and the second-order Zeeman effect is taken into account during the holding time of the lower arm, as well as the effect of the second-order Zeeman over the whole sequence.
@@ VII.8.1 Details of the numerical simulation
The free parameters of the simulation are the bias field in the x and y directions and the exact value of the collimation current. When these parameters are varied under the constraint that the lower arm trajectory is flat and at the correct distance from the chip surface, the interferometer phase varies within a range of ±$\pm 0.07$ rad at the longest holding time. Note that these parameters were not used for fitting the phase but rather to fit other parameters of the experiment mentioned above; thus the numerical simulation curve in Fig. 3 can be considered as a curve without fitting parameters related to the phase.
@@
VII.8.2 Effect of wave packet shape on the phase
The numerical simulation includes the effect of the wave-packet shape, which evolves due to single particle effects (internal kinetic energy at the wave-packet center and response to the curved potentials) and atom-atom interactions. To estimate the effects of the wave-packet shape on the phase, we compare the total phase calculated from the overlap integral to the phase of the classical trajectory of the COM. The result is shown in Fig. 9. The total phase accumulation in the experiment amounts to $80\rm\,rad$ while the contribution due to the shape of the wave packet is in the range of 00 to $1\rm\,rad$. The relative contribution of the “non trajectory” phase is as high as $7\%$ for short interferometer durations, but reduces for longer interferometer times. The exact contribution due to the shape of the wave packet depends on the exact shape of the wave packet and the specific experimental conditions, and can be one of the sources of the discrepancy between the numerical simulation and the experimental data. The numerical simulation also includes the effect of the ambient gradient in the chamber, which was measured to be $0.7\,\rm G/cm$ (equivalent to an acceleration of $0.2\,\rm m/s^{2}$ along gravity of the $|1\rangle$ state, see Sec. L for more details) and the idle current of $0.55\,\rm mA$ we apply to counteract the effect of the ambient gradient. The shape of the pulses of current in the SG wires is also included in the simulation. A detailed description of the pulses is given in Sec. J.
Figure 9: The different phase contributions calculated in the numerical simulation. (A) The total phase (blue solid line), was calculated from the overlap integral between the two wave packets at the end of the interferometer (the timing of the last $\pi/2$ pulse). This total phase takes into account effects of wave packet shape and atom-atom interaction. The red dashed line is the “path phase”, the phase accumulation due to the classical trajectory of the COM. (B) The difference between the two calculations, shows that the “non trajectory” contributions range between $-1.1$ to $0.4\rm\,rad$, out of the total $80\rm\,rad$ of the interferometer. (C) The relative part of the “non trajectory” is as high as $15\%$ for short interferometer duration, but reduces to less than $2\%$ for durations longer than $1.2\rm\,ms$.
VII.9 Data analysis for figures 2 and 3

### v4

@@ G.2 Derivation of the phase difference using the action
$\int_{t_{i}}^{t_{f}}a_{1}(t)dt=\int_{t_{i}}^{t_{f}}a_{2}(t)dt\,$ (S11)
which, in this case, gives
$a_{kick}\cdot 2T_{kick}=a_{hold}\cdot T_{h}\,.$ (S12)
Setting $a_{hold}=g$ confirms that the reference path is stationary in the laboratory frame. Performing the integrals, we get
$\phi_{1}=\frac{mg^{2}}{3\hbar}\left(-T^{3}-T^{2}T_{kick}-TT_{kick}^{2}-TT_{kick}T_{d}+2T_{kick}^{3}+7T_{kick}^{2}T_{d}+8T_{kick}T_{d}^{2}+3T_{d}^{3}\right)$ (S13)
$\phi_{2}=\frac{2mg^{2}}{3\hbar}\left(T_{kick}+T_{d}\right)^{3}$ (S14)
$\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}+T^{2}T_{kick}+T\left(T_{kick}^{2}+T_{kick}T_{d}\right)-T_{d}\left(T_{kick}+T_{d}\right)^{2}\right).$ (S15)
Taking the limit of a short kick we get
$\lim_{T_{kick}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}-T_{d}^{3}\right),$ (S16)
and the limit of short delay we get
$\lim_{T_{kick},T_{d}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}T^{3}.$ (S17)
Eq.S17 represents the Gedanken experiment in which the beam splitter is a delta pulse such that $T_{kick},T_{d}\to 0$.
The result of the analytical model given in Eq. (S15) (and in Eq. (4) of the main text) is plotted in Fig. 3 of the main text using the experimental values of $T_{kick}=80\rm\,\mu s$ and $T_{d}=71+6\rm\,\mu s$ where we include half the rise time of the holding pulse $\tau_{\rm hold}/2=6\rm\,\mu s$ in the delay. We also include the contribution to the phase from the force on the state $|0\rangle$ due to the second-order Zeeman (SOZ) effect. This force is opposite to the force on the $|1\rangle$ state during the gradient pulses, and its value is about 1% of the latter. During the holding pulse, the acceleration of the state $|0\rangle$ due to SOZ is calculated to be ${a}_{SOZ}=0.10\,\rm m/s^{2}$ (see Sec. K) for details). We include this effect in the analytical prediction for the phase (Eq. 17) by replacing the gravitational acceleration $g$ by $g_{\rm eff}=g+a_{\rm SOZ}=9.91\,\rm m/s^{2}$. This estimation agrees with a model in which $a_{SOZ}$ is explicitly included in the calculation to within $0.03\,\rm rad$.
G.3 Calculating the actions in both frames
Next, to understand the symmetry of the QGI, as emphasized in the introduction, we calculate for the Gedanken experiment the actions associated with the classical trajectories of the reference and the ballistic WPs as seen from both frames of reference, keeping track of the contributions from the kinetic and potential energies as well as the pulses, which we model for the sake of simplicity by delta functions.
We start by discussing the situation in the Newtonian frame. Here the reference WP remains at rest for all times, since the gravitational acceleration $g$ is compensated by the acceleration $a$ provided by the magnetic field gradient. Hence, the contributions due to the kinetic and potential energies vanish. Moreover, there are no pulses. Thus, the total action of the reference WP vanishes as indicated in Table S1.
In contrast, two parts contribute to the total action of the ballistic WP. Due to the non-vanishing momentum $mv_{0}$ of the atom originating from the first pulse, the accelerated motion in the linear gravitational potential provides us with non-vanishing actions due to the kinetic as well as the potential energies. Their structure shown in the second row of Table S1 becomes apparent when we recall that the velocity of this motion consists of the constant velocity $v_{0}$ and the one caused by the gravitational acceleration $g$. The action due to this kinetic energy involves the integral of the square of this time-dependent velocity creating the sum of three terms: (ii) The kinetic energy $mv_{0}^{2}/2$ multiplied by the interferometer time $2T$. (iiii) A cross term between the velocity $v_{0}$ and the acceleration $g$ and (iiiiii) the square of the acceleration times $(2T)^{3}$.
This scaling in $v_{0}g$ and $g^{2}$ is crucial. Indeed, when we replace $v_{0}$ by $v_{0}$ and $g$ by a-a as defined by the Einsteinian frame discussed in the second column of Table S1, the expressions, now occurring for the reference rather than the ballistic WP, are identical.
Next we turn to the contribution from the potential energy of the ballistic WP in the Newtonian frame. Since the ballistic WP starts from the origin of the coordinate axis, the motion is a combination of the linear motion due to $v_{0}$, and the gravitational acceleration $g$ giving rise to two terms in the action which are identical in size and sign to the last two terms of the action due to the kinetic energy. Hence, these terms also satisfy the symmetry in $v_{0}g$ and $g^{2}$.
It is interesting that in the Newtonian frame the delta function pulses do not contribute to the actions of the ballistic WP in the Newtonian frame, and this for two reasons: First, the corresponding potential energy is proportional to the coordinate at the time of the pulse and second, the ballistic WP starts at $z=0$ and concludes its journey at $z=0$.
As a result, the total action of the ballistic WP is given by the ones of the kinetic and potential energies which due to the closing condition connecting $v_{0}$ and $T$ allows us to combine the three terms in a single one shown in the Table S1, determining the difference of the actions corresponding to the two WPs.
Finally, we address the Einsteinian frame which is attached to the motion of the ballistic WP. As mentioned before, the two WPs exchange their roles compared to the Newtonian frame. Hence, now the action of the ballistic WP vanishes but the one of the reference WP is non-vanishing arising from the magnetic field gradient in the same way as for the ballistic WP in the Newtonian frame.
Moreover, the symmetry of the terms discussed before, ensures that the origin of this contribution is identical to one of the terms of the ballistic WP in the Newtonian frame.
Since we consider the difference of the actions of the two WPs there must be a sign difference between the two frames. Indeed, in the Newtonian frame the action of the reference WP vanishes, whereas for the Einsteinian frame it is that of the ballistic WP.
G.4 Derivation of the transformation phase between the Newtonian and Einsteinian frames using Galilean transformation
Here we re-derive the transformation phase between the Newtonian and Einsteinian frames using the Galilean transformation without relying in any way on the EP. For this purpose, we do not assume equality of the inertial and gravitational masses and use the Newton equation $m_{i}\ddot{z}_{cl}=-m_{g}g$, to connect
$\psi_{N}(z,t)=e^{i\chi(z,t)}\psi_{E}\left(z-v_{0}t+\frac{1}{2}\frac{m_{g}}{m_{i}}gt^{2},t\right),\;\text{where}\;\;\;\chi(z,t)\equiv\chi_{0}(t)+\chi_{1}(t)z$ (S18)
the time-dependent wave function $\psi_{N}$ in the Newtonian frame and the one $\psi_{E}$ in the Einsteinian frame, that is the one moving along the classical trajectory $z_{cl}(t)=v_{0}t-(1/2)(m_{g}/m_{i})gt^{2}$ with a initial velocity $v_{0}$. For writing the phase $\chi(x,t)$ as the linear function of $z$ with two unknown functions of time $\chi_{0}(t)$ and $\chi_{1}(t)$, we have used the fact that the particle always increases its momentum in a linear potential.
To derive the analytical forms of $\chi_{0}(t)$ and $\chi_{1}(t)$, one has to insert the function $\psi_{N}(z,t)$ into the Schrödinger equation
$i\hbar\frac{\partial}{\partial t}\psi_{N}(z,t)=\left(-\frac{\hbar^{2}}{2m_{i}}\frac{\partial^{2}}{\partial z^{2}}+m_{g}gz\right)\psi_{N}(z,t)$ (S19)
in the Newtonian frame and use then the Schrödinger equation
$i\hbar\frac{\partial}{\partial t}\psi_{E}(z^{\prime},t)=-\frac{\hbar^{2}}{2m_{i}}\frac{\partial^{2}}{\partial z^{\prime 2}}\psi_{E}(z^{\prime},t)$ (S20)
in the Einsteinian frame, the particle is at rest in this frame but undergoes free quantum dynamics.
As a result, we arrive at
$\begin{split}\chi_{0}(t)=&-\frac{1}{\hbar}\int_{0}^{t}d\tau\frac{\left(m_{i}v_{0}-m_{g}g\tau\right)^{2}}{2m_{i}}=-\frac{1}{2\hbar m_{i}}\left(m_{i}^{2}v_{0}^{2}t-m_{i}m_{g}v_{0}gt^{2}+\frac{1}{3}m_{g}^{2}g^{2}t^{3}\right)\\ \chi_{1}(t)=&\frac{m_{i}v_{0}-m_{g}gt}{\hbar}.\end{split}$ (S21)
In the case of $v_{0}=0$ and $m_{i}=m_{g}$, Eq. (S18) gives Eq. (1) of the main text.
Finally, we obtain the phase difference measured by the QGI given by Eq. (2) of the main text, using only the Galilean transformation, Eq. (S18). To do it, we consider the wave function $\psi_{N}$ at $t=2T$ and use the closing condition $T=m_{i}v_{0}/(m_{g}g)$. As a result, the spatial argument in $\psi_{E}$ reduces to $z$, corresponding to perfect overlap of the reference and ballistic WPs in the coordinate space. Moreover, the second pulse employed at $t=2T$ imprints again the momentum $p_{0}=m_{i}v_{0}$ onto the wave function and thus cancels the position-dependent phase $\chi_{1}(2T)z=-m_{i}v_{0}z/\hbar$. Hence, the QGI is closed in both momentum and position space. The phase difference therefore coincides with the phase between $\psi_{N}(z,2T)$ and $\psi_{E}(z,2T)$, that is
$\psi_{N}(z,2T)=e^{i\phi_{ballistic}}\psi_{E}\left(z,2T\right),$ (S22)
where the ballistic phase is entirely determined by $\chi_{0}(2T)$ and reads
$\phi_{ballistic}=\chi_{0}(2T)=-\frac{1}{2\hbar m_{i}}\left(\frac{8}{3}m_{g}^{2}g^{2}T^{3}-2m_{i}^{2}v_{0}^{2}T\right)=-\frac{1}{2\hbar m_{i}}\left(\frac{8}{3}-2\right)m_{g}^{2}g_{2}T^{3}=-\frac{m_{g}^{2}}{3\hbar m_{i}}g^{2}T^{3}.$ (S23)
Here we have used the closing condition, $T=m_{i}v_{0}/(m_{g}g)$.
Table S1: Contributions to the actions of the center-of-mass motion of the reference and the ballistic WPs in the QGI viewed from the Newtonian and the Einsteinian frames. Note that the closing conditions to obtain the respective total actions are written in round brackets whereas the levitation condition for the respective difference in actions are given by the expression in square brackets. In the Newtonian frame the reference WP is levitated and remains at rest at the origin. The Einsteinian frame is attached to the ballistic WP. Thus, in this frame the WP remains at rest at the origin yielding only vanishing actions. This is also the case for the reference WP but in the Newtonian frame.
actions frame Newtonian Einsteinian
reference WP kinetic 0 $mv_{0}^{2}T-2mv_{0}aT^{2}+\frac{4}{3}ma^{2}T^{3}$
potential 0 $2mv_{0}aT^{2}+\frac{4}{3}ma^{2}T^{3}$
pulses no pulses 0
total 0 $\frac{1}{3}ma^{2}T^{3}$ $(v_{0}=aT)$
ballistic WP kinetic $mv_{0}^{2}T-2mv_{0}gT^{2}+\frac{4}{3}mg^{2}T^{3}$ 0
potential $2mv_{0}gT^{2}+\frac{4}{3}mg^{2}T^{3}$ 0
pulses 0 no pulses
total $\frac{1}{3}mg^{2}T^{3}$ ($v_{0}=gT$) 0
difference in total actions $\frac{1}{3}mg^{2}T^{3}$ $\frac{1}{3}ma^{2}T^{3}$ [a=g][a=-g]
H Numerical simulation
H.1 Details of the numerical simulation
The numerical simulation calculates the dynamics of the atomic BEC during the trapping, release, DKC, and interferometer sequence. The simulation is based on the model of WP evolution given in [53], which is extended to include the effects of rotations. To use the model with curved magnetic potentials, we locally approximate the magnetic potential by a quadratic potential and calculate the evolution of each WP in the potential. The simulation predicts the output of the SGI by calculating the overlap integral of the two final WPs, resulting in a complex value, where the visibility is the magnitude of this integral and the phase is the argument.
@@ H.1 Details of the numerical simulation
[…]
@@
The free parameters of the simulation are the bias field in the x and y directions and the exact value of the collimation current. When these parameters are varied under the constraint that the lower arm trajectory is flat and at the correct distance from the chip surface, the interferometer phase varies within a range of $\pm 0.07$ rad at the longest holding time. Note that these parameters were not used for fitting the phase but rather to fit other parameters of the experiment mentioned above; thus the numerical simulation curve in Fig. 3 can be considered as a curve without fitting parameters related to the phase.
@@
H.2 Effect of WP shape on the phase
The numerical simulation includes the effect of the wave-packet shape, which evolves due to single particle effects (internal kinetic energy at the wave-packet center and response to the curved potentials) and atom-atom interactions. To estimate the effects of the wave-packet shape on the phase, we compare the total phase calculated from the overlap integral to the phase of the classical trajectory of the COM. The result is shown in Fig. S4. The total phase accumulation in the experiment amounts to $80\rm\,rad$ while the contribution due to the shape of the WP is in the range of 0 to $1\rm\,rad$. The relative contribution of the “non trajectory” phase is as high as $7\%$ for short interferometer durations, but reduces for longer interferometer times. The exact contribution due to the shape of the WP depends on the exact shape of the WP and the specific experimental conditions, and can be one of the sources of the discrepancy between the numerical simulation and the experimental data. The numerical simulation also includes the effect of the ambient gradient in the chamber, which was measured to be $0.7\,\rm G/cm$ (equivalent to an acceleration of $0.2\,\rm m/s^{2}$ along gravity of the $|1\rangle$ state, see Sec. L for more details) and the idle current of $0.55\,\rm mA$ we apply to counteract the effect of the ambient gradient. The shape of the pulses of current in the SG wires is also included in the simulation. A detailed description of the pulses is given in Sec. J.
Figure S4: The different phase contributions calculated in the numerical simulation. (A) The total phase (blue solid line), was calculated from the overlap integral between the two WPs at the end of the interferometer (the timing of the last $\pi/2$ pulse). This total phase takes into account effects of WP shape and atom-atom interaction. The red dashed line is the “path phase”, the phase accumulation due to the classical trajectory of the COM. (B) The difference between the two calculations, shows that the “non trajectory” contributions range between 1.1-1.1 to $0.4\rm\,rad$, out of the total $80\rm\,rad$ of the interferometer. (C) The relative part of the “non trajectory” is as high as $15\%$ for short interferometer duration, but reduces to less than $2\%$ for durations longer than $1.2\rm\,ms$.
I Data analysis for figures 2 and 3